# Feature-Based Interpretable Optimization

### v1

@@ 1.3 Contribution and Structure
In this paper, we introduce a new, feature-based framework for interpretable optimization. While a previous approach proposed in [GH23] maps instance parameters to solutions with a decision tree, we propose to map instance features to solution features instead. Given solution features can be regarded as a meta-solution, i.e., a representative for a set of feasible solutions for the optimization problem. We therefore extend the flexibility of the setting, thus resulting in better solutions for the decision-maker. This extension brings several further benefits: (i) It is easier to explain a meta-solution as we only need to describe the key features corresponding to the solution set instead of an individual solution. The features are of a smaller dimension and thus are easier to explain. (ii) Using general instance features, rather than only working with coefficients in the objective function, expands the range of problems the framework can cope with. In particular, by using instance and solution features, the dimension of the underlying problem no longer needs to be fixed. The domain now can depend on the instance features and hence, the framework can be applied to more general settings. (iii) By providing meta-solutions, we assume that a user is able to identify the specifics of the preferred solution, giving credit to his or her domain knowledge. Additionally, but somewhat contrarily, the optimization rule can be provided to a stakeholder as an argument for why the current solution was implemented. (iv) From a managerial perspective, providing a meta-solution rather than a specific solution allows the planner greater flexibility in demonstrating that the implemented solution aligns with pre-established rules. However, this approach somewhat calls for ensuring that meta-solutions remain as general as possible to accommodate various interpretations and implementations.
@@
The remainder of this paper is structured as follows. In Section 2, we formally define our framework for feature-based interpretable optimization and outline how this framework can be applied in the context of different example problems. We then derive mixed-integer programming models in Section 3, focussing on the knapsack problem and the shortest path problem, where we also provide hardness proofs. In Section 4, we introduce heuristics for our framework. Our methods are applied in Section 5, where we present numerical experiments involving artificial and real-world data. Finally, Section 6 summarizes our paper and provides further research questions.
@@
Throughout the paper, we focus on univariate, binary decision trees of fixed depth to ensure comprehensibility. Hence, by performing easy queries on the instance features to obtain a meta-solution, the optimization process becomes interpretable. Other interpretable mappings from instance features to solution features could be used instead. We use the notation $[n]:=\{1,\ldots,n\}$ to denote index sets and write vectors in bold font. Furthermore, we use “$\min$” and “$\max$” in optimization problems rather than “$\inf$” and “$\sup$” to denote the direction of optimization, and assume that problems are sufficiently well-posed to allow for the existence of the minimum or maximum, respectively.
2 Framework for Interpretable Optimization using Feature-Based Meta-Solutions
@@ 2.1 The Framework
where $X\subseteq\mathbb{R}^{n}$ denotes the set of feasible solutions, $n\in\mathbb{N}$ is the number of variables, and $c:X\rightarrow\mathbb{R}$ is an objective function. We assume that an overarching problem domain ${\mathcal{X}}$ can be modeled, which dictates specific constraints but might lack a fixed dimension. For example, ${\mathcal{X}}$ may denote the union of all paths or matchings within graphs of various sizes. In this context, while the fundamental problem setting ${\mathcal{X}}$ remains the same, several parameters, referred to as instance features, are subject to variation. Let $\Phi_{I}\subseteq\mathbb{R}^{p}$ be the space of possible realizations of the $p\in\mathbb{N}$ instance features. For $\phi_{I}\in\Phi_{I}$ the optimization problem is given by
$\min_{{\bm{x}}\in{\mathcal{X}}(\phi_{I})}c(\phi_{I},{\bm{x}}),$
@@
with ${\mathcal{X}}(\phi_{I})\subseteq{\mathcal{X}}$ and $c:\Phi_{I}\times{\mathcal{X}}\rightarrow\mathbb{R}$, i.e., the instance features can affect the solution space as well as the objective function. We assume a given finite set of $N\in\mathbb{N}$ candidate scenarios also referred to as training scenarios, each represented by its respective instance features $\phi_{I}^{j}\in\Phi_{I}$, $j\in[N]$, with probabilities $\bm{p}\in[0,1]^{N}$, $\sum_{j\in[N]}p_{j}=1$. We want to find an interpretable optimization rule, that maps a scenario given by its instance features to a meta-solution, i.e., a collection of solutions characterized by easily understandable features shared among them. To that end, we assume there exists a set of $F\in\mathbb{N}$ solution features (stemming from problem-specific knowledge) that describe a solution in an easily comprehensible manner. Let $\Phi_{S}\subseteq\mathbb{R}^{F}$ be the solution feature space and for any solution ${\bm{x}}\in{\mathcal{X}}$, let $\phi_{S}({\bm{x}})\in\Phi_{S}$ extract solution features from solution ${\bm{x}}$.
@@
The features are selected carefully to i) provide meaningful insights into the structure of the solution, ii) be easily comprehensible for the stakeholder, and iii) allow a practitioner to easily compute a solution when having these features at hand. When both the provided features are comprehensible and the circumstances dictating the employment of specific features are elucidated, a more interpretable solution process is established. Additionally, in contrast to [GH23], by relying on the instance and solution features, the problem dimensions can change, while the provided optimization rule still is intact.
@@
We aim to find an easily comprehensible optimization rule $a:\Phi_{I}\rightarrow\Phi_{S}$ that maps new instances (given by its instance features) to a set of feasible solutions (given by solution features). In particular, it describes the characteristics of a good solution, by providing information on the characteristics of each solution feature, depending on the instance features. Solution features do not uniquely describe a solution, but rather a set of solutions. We call $\phi_{S}\in\Phi_{S}$ a meta-solution, i.e., the solution space specified by the feature characteristics. The used solution in a new instance then ought to conform to the specified structure of the meta-solution, while still giving the decision maker some leeway. Hence, even if two different instances call for the same meta-solution, the implemented solutions can differ. We sometimes refer to specific solutions $\bm{x}$ as “micro”-solutions to emphasize the difference.
@@
As the main goal is to obtain a comprehensible optimization rule, the structure of allowed optimization rules $A\subseteq\{a:\Phi_{I}\rightarrow\Phi_{S}\}$ is crucial. In this paper, we focus on decision trees with a restricted depth, where the branching decisions outline the rule. This way, the number of meta-solutions that can occur is small and the mapping of a scenario to a meta-solution is easily comprehensible. Note, however, that other approaches, e.g., linear optimization rules, also can be employed in this framework.
@@
$\min_{a\in A}\mu_{\bm{p}}\left(C(\phi_{I}^{1},a(\phi_{I}^{1})),\ldots,C(\phi_{I}^{N},a(\phi_{I}^{N}))\right),$
@@
$C(\phi_{I},a(\phi_{I}))=\min_{\begin{subarray}{c}\bm{x}\in{\mathcal{X}}(\phi_{I})\\ \phi_{S}(\bm{x})=a(\phi_{I})\end{subarray}}c(\phi_{I},\bm{x}),$
i.e., after the optimization rule has outlined a meta-solution $a(\phi_{I})$, the best solution within the solution space is selected that exhibits these features.
The general decision criterion $\mu_{\bm{p}}$ allows for the appropriate handling of various settings, e.g., for the Laplace criterion our optimization problem is given by
$\min_{a\in A}\sum_{j\in[N]}\min_{\begin{subarray}{c}\bm{x}\in{\mathcal{X}}(\phi_{I}^{j})\\ \phi_{S}(\bm{x})=a(\phi_{I}^{j})\end{subarray}}c(\phi_{I}^{j},\bm{x})\ ,$
@@
$\min_{a\in A}\max_{j\in[N]}\min_{\begin{subarray}{c}\bm{x}\in{\mathcal{X}}(\phi_{I}^{j})\\ \phi_{S}(\bm{x})=a(\phi_{I}^{j})\end{subarray}}c(\phi_{I}^{j},\bm{x})\ .$
@@ 2.2 Scope of the Framework
Let us first walk through the progressive expansion of the knapsack problem and how our framework can be applied to it. Consider a combinatorial knapsack problem with $n$ items and a budget $C$. Given a profit vector $\bm{c}$ and a weight vector $\bm{w}$ we can identify an optimal knapsack solution using the following optimization problem:
@@
We now consider a situation with $N$ profit scenarios, that is, we assume that the instance feature vector $\phi_{I}$ corresponds to a vector of item profits. Let these be indexed by $j$. To leverage our framework, we assign each item to one of $F$ different categories. Let $I_{f}$ denote the set of items assigned to category $f\in[F]$. The meta-solution $\phi_{S}$ then consists of specifying the budget for each category, i.e., the $F$ solution features specify the amount $C_{f}$ of the budget spent on category $f\in[F]$. We can write the optimization problem for calculating such a budget as
@@
$C_{f}\geq 0$ $\forall f\in[F].$
In the above setting, we only obtain one single meta-solution (budget capacities for each category) to cover all scenarios. Note that the specific realization of the solution still has some leeway, as the specific selection of items is not prescribed by the meta-solution: it only tells us to use a budget of $C_{f}$ for items of category $f$ but does not prescribe which exact items to pick. An even more flexible alternative would be to have multiple possible meta-solutions (though less than the number of scenarios) and choose one among them for any given scenario. We index this set of meta-solutions by $k\in[K]$. We can then write the problem as
@@
s.t. $\sum_{f\in[F]}C_{f}^{k}\leq C$ $\forall k\in[K]$
$\sum_{i\in I_{f}}w_{i}x_{ji}\leq C_{f}^{k}+M(1-\ell^{k}_{j})$ $\forall f\in[F],\ j\in[N],\ \forall k\in[K]$
@@
$C^{k}_{f}\geq 0$ $\forall f\in[F],\ k\in[K]\,,$
where the goal is to find a function $f(\bm{c}_{j},\bm{\theta})$ that always takes the form of a decision tree. The splits $\bm{\theta}$ must be optimized so that the function maps instance features (the profit vector) to a meta-solution through this decision tree while maximizing the average profit.
We can understand the final knapsack problem in terms of the elements introduced in Section 2.1. Here, the set $\mathcal{X}$ corresponds to the budget constraint of the knapsack problem, which in this example is not affected by the instance features. However, the constraints forming the set are modified to model the meta-solutions. Each instance feature $\phi_{I}^{j}$ reflects a different cost vector for the objective. The knapsack problem as modeled is equivalent to assigning equal probabilities to all scenarios. The solution features $\phi_{S}$ thus correspond to a meta-solution, which is equivalent to a capacity vector $(C_{1}^{k},\dots,C_{F}^{k})$ with $\Phi_{S}\subseteq\mathbb{R}^{F}$. Given a meta-solution and instance features, the function $C(\phi_{I},\phi_{S})$ from our framework calculates the optimal objective value obtained from selecting the best singular solution obeying the meta-solution on that instance. The set $A$ is the space of all decision trees that map instance features to a meta-solution. We parameterize this set by $\bm{\theta}$. Finally, $\mu_{\bm{p}}$ represents how the objective values of different instances are aggregated which is by summing them together, which is equivalent to the Laplace criteria.
@@
Assume we are given a graph $G=(V,E)$ representing a road network and nodes $s,t\in V$ representing a starting point and a destination. Furthermore, let us assume that traffic follows a pattern as shown in Figure 1, which depicts the day of the week on the horizontal axis, and the time of day on the vertical axis. We assume that there are four main traffic scenarios that can occur at different times of the day throughout the week. We want to provide a comprehensible optimization rule for selecting a route, depending on the time of day and the day of the week.
Figure 1: Traffic scenarios.
To make our example more specific, we are using the graph from [CDG19], representing Chicago’s street network, which we have divided into eleven districts (see gray boxes in Figure 2).
@@
The edges are colored if travel speed on them is slower than usual. The colors range from yellow (slightly reduced) to dark red (significantly slower) and provide information about the intensity of the speed reduction.
For the task of traveling from the northwest to the southeast of the city, let us first consider the original approach of [GH23].
@@
Figure 3: Optimization rule obtained from using the framework presented in [GH23].
Figure 3 illustrates the optimization rule through a binary decision tree with a maximum depth of two. In each of the four scenarios identified by the decision tree’s splits, a specific $s$-$t$ path must be determined.
However, this approach may be too restrictive and diverges from how directions are typically explained in real life, where one would emphasize prominent landmarks, districts, streets, or intersections along the route, rather than detailing every road segment to be traveled. Therefore, we want to provide an optimization rule that uses solution features. In Figure 4, another binary decision tree is provided which functions as an optimization rule. However, instead of providing $s$-$t$ paths in the leaves, we now provide meta-solutions (shown in Figure 5), i.e., we provide features that any solution should possess in the given scenario. In this specific case, we select as solution features, the sequence of city districts that one should traverse to reach the destination.
@@
(b) Meta-solution 2 and 3.
(c) Meta-solution 4.
Figure 5: Meta-solutions for the new framework.
Districts shaded in gray have to be visited as per the order indicated by the bold gray arrows. Unshaded districts may not be used. This way, we obtain a more comprehensible specification (specify a few features, rather than a lot of edges) and give the user more leeway in the implementation. In particular, any solution that follows the specified path through the districts is suitable.

### v3

@@ 1.3 Contribution and Structure
In this paper, we address the research question of how to obtain interpretable surrogates for optimization and introduce a novel feature-based framework for deriving such optimization rules. While the solution-based approach seeks to find surrogates that map cost parameters to solutions, we instead propose mapping instance features to solution features. An instance feature can include not only parameters of the given optimization problem but also contextual data. Given solution features can be regarded as a meta-solution, i.e., a representative for a set of feasible solutions for the optimization problem. We therefore extend the flexibility of the setting, thus resulting in better solutions for the decision-maker. We also give insights into the question of how high the cost of enforcing interpretability in optimization is through a computational study.
@@
Our feature-based framework brings several further impacts: (i) It is easier to explain a meta-solution as we only need to describe the key features corresponding to the solution set instead of an individual solution. Furthermore, features are of a smaller dimension and thus are easier to describe. This will strengthen the acceptance of managerial decisions. (ii) Using general instance features, rather than only working with coefficients in the objective function, expands the range of problems the framework can cope with. In particular, by using instance and solution features, the dimension of the underlying problem no longer needs to be fixed. The domain now can depend on the instance features, and hence, the framework can be applied to more general settings. (iii) By providing meta-solutions, we assume that a user is able to identify the specifics of the preferred solution, giving credit to his or her domain knowledge. Additionally, but somewhat contrarily, the optimization rule can be provided to a stakeholder as an argument for why the current solution was implemented. (iv) From a managerial perspective, providing a meta-solution rather than a specific solution allows the planner greater flexibility in demonstrating that the implemented solution aligns with pre-established rules. However, this approach somewhat calls for ensuring that meta-solutions remain as general as possible to accommodate various interpretations and implementations.
@@
The remainder of this paper is structured as follows. In Section 2, we formally define our framework for feature-based interpretable surrogates and outline how this framework can be applied in the context of different example problems. We then derive mixed-integer programming models in Section 3, focusing on the knapsack problem and the shortest path problem, where we also provide hardness proofs. In Section 4, we introduce heuristics for our framework. Our methods are applied in Section 5, where we present numerical experiments involving artificial and real-world data. Finally, Section 6 summarizes our paper and provides further research questions.
@@
Throughout the paper, we focus on univariate, binary decision trees of fixed depth to ensure comprehensibility. Hence, by performing easy queries on the instance features to obtain a meta-solution, the optimization process becomes interpretable. Other interpretable mappings from instance features to solution features could be used instead. We use the notation $n]:=\{1,\ldots,n\}$ to denote index sets and write vectors in bold font. Furthermore, we use $\min$” and $\max$” in optimization problems rather than $\inf$” and $\sup$” to denote the direction of optimization and assume that problems are sufficiently well-posed to allow for the existence of the minimum or maximum, respectively.
2 Feature-Based Framework for Finding Interpretable Surrogates
@@ 2.1 The Framework
where $X\subseteq\mathbb{R}^{n}$ denotes the set of feasible solutions, $n\in\mathbb{N}$ is the number of variables, and $c:X\rightarrow\mathbb{R}$ is an objective function. We assume that an overarching problem domain $\mathcal{X}$} can be modeled, which dictates specific constraints but might lack a fixed dimension. For example, $\mathcal{X}$} may denote the union of all paths or matchings within graphs of various sizes. In this context, while the fundamental problem setting $\mathcal{X}$} remains the same, several parameters, referred to as instance features, are subject to variation. Let $\Phi_{I}\subseteq\mathbb{R}^{F_{I}}$ be the space of possible realizations of the $F_{I}\in\mathbb{N}$ instance features. For $\bm{\phi}_{I}\in\Phi_{I}$ the optimization problem is given by
$\min_{{\bm{x}}\in{\mathcal{X}}(\bm{\phi}_{I})}c(\bm{\phi}_{I},{\bm{x}}),$
@@
with ${\mathcal{X}}(\bm{\phi}_{I})\subseteq{\mathcal{X}}$ and $c:\Phi_{I}\times{\mathcal{X}}\rightarrow\mathbb{R}$, i.e., the instance features can affect the solution space as well as the objective function. We assume a given finite set of $N\in\mathbb{N}$ candidate scenarios also referred to as training scenarios, each represented by its respective instance features $\bm{\phi}_{I}^{j}\in\Phi_{I}$, $j\in[N]$, with probabilities $\bm{p}\in[0,1]^{N}$, $\sum_{j\in[N]}p_{j}=1$. We want to find an interpretable optimization rule. This is a function that maps a scenario given by its instance features to a meta-solution, i.e., a collection of solutions characterized by easily understandable features shared among them. To that end, we assume there exists a set of $F_{S}\in\mathbb{N}$ solution features (stemming from problem-specific knowledge) that describe a solution in an easily comprehensible manner. Let $\Phi_{S}\subseteq\mathbb{R}^{F_{S}}$ be the solution feature space and for any solution ${\bm{x}}\in{\mathcal{X}}$, let $\bm{\phi}_{S}({\bm{x}})\in\Phi_{S}$ extract solution features from solution $\bm{x}$}.
@@
The features are selected carefully to i) provide meaningful insights into the structure of the solution, ii) be easily comprehensible for the stakeholder, and iii) allow a practitioner to easily compute a solution when having these features at hand. When both the provided features are comprehensible and the circumstances dictating the employment of specific solution features are elucidated, an interpretable surrogate for the optimization process is established. Additionally, in contrast to solution-based interpretable surrogates, by relying on the instance and solution features, the problem dimensions can change, while the provided optimization rule remains intact.
@@
We aim to find an easily comprehensible optimization rule $a:\Phi_{I}\rightarrow\Phi_{S}$ that maps new instances (given by their instance features) to a set of feasible solutions (given by solution features). In particular, it describes the characteristics of a good solution by providing information on the specifics of each solution feature, depending on the instance features. Solution features do not uniquely describe a solution, but rather a set of solutions. We call $\bm{\phi}_{S}\in\Phi_{S}$ a meta-solution, i.e., the solution space specified by the feature characteristics. The used solution in a new instance ought to conform to the specified structure of the meta-solution, while still giving the decision maker some leeway. Hence, even if two different instances call for the same meta-solution, the implemented solutions can differ. We sometimes refer to specific solutions $\bm{x}$ as “micro”-solutions to emphasize the difference.
@@
As the main goal is to obtain a comprehensible optimization rule, the structure of allowed optimization rules $A\subseteq\{a:\Phi_{I}\rightarrow\Phi_{S}\}$ is crucial. In this paper, we focus on decision trees with a restricted depth, where the branching decisions outline the rule. This way, the number of meta-solutions that can occur is small and the mapping of a scenario to a meta-solution is easily comprehensible. Note, however, that other approaches, e.g., linear optimization rules, can also be employed in this framework.
@@
$\min_{a\in A}\mu_{\bm{p}}\left(C(\bm{\phi}_{I}^{1},a(\bm{\phi}_{I}^{1})),\ldots,C(\bm{\phi}_{I}^{N},a(\bm{\phi}_{I}^{N}))\right),$
@@
$C(\bm{\phi}_{I},a(\bm{\phi}_{I}))=\min_{\begin{subarray}{c}\bm{x}\in{\mathcal{X}}(\bm{\phi}_{I})\\ \bm{\phi}_{S}(\bm{x})=a(\bm{\phi}_{I})\end{subarray}}c(\bm{\phi}_{I},\bm{x}),$
i.e., after the optimization rule has outlined a meta-solution $a(\bm{\phi}_{I})$, the best solution within the solution space is selected that exhibits these features.
The general decision criterion $\mu_{\bm{p}}$ allows for the appropriate handling of various settings, e.g., for the Laplace criterion, our optimization problem is given by
$\min_{a\in A}\sum_{j\in[N]}\min_{\begin{subarray}{c}\bm{x}\in{\mathcal{X}}(\bm{\phi}_{I}^{j})\\ \bm{\phi}_{S}(\bm{x})=a(\bm{\phi}_{I}^{j})\end{subarray}}c(\bm{\phi}_{I}^{j},\bm{x})\ ,$
@@
$\min_{a\in A}\max_{j\in[N]}\min_{\begin{subarray}{c}\bm{x}\in{\mathcal{X}}(\bm{\phi}_{I}^{j})\\ \bm{\phi}_{S}(\bm{x})=a(\bm{\phi}_{I}^{j})\end{subarray}}c(\bm{\phi}_{I}^{j},\bm{x})\ .$
Please note that the framework is designed to be highly general and applicable to a wide range of optimization problems. However, in the following sections, we focus on combinatorial optimization problems.
@@ 2.2 Scope of the Framework
Let us first walk through the progressive expansion of the knapsack problem and how our framework can be applied to it. Consider a combinatorial knapsack problem with $n$ items and a budget $C$. Given a profit vector $\bm{c}$ and a weight vector $\bm{w}$, we can identify an optimal knapsack solution using the following optimization problem:
@@
We now consider a situation with $N$ profit scenarios, that is, we assume that the instance feature vector $\bm{\phi}_{I}$ corresponds to a vector of item profits. Let these be indexed by $j$. To leverage our framework, we assign each item to one of $F_{S}$ different categories. Let $I_{f}$ denote the set of items assigned to category $f\in[F_{S}]$. The meta-solution $\bm{\phi}_{S}$ then consists of specifying the budget for each category, i.e., the $F_{S}$ solution features specify the amount $C_{f}$ of the budget spent on category $f\in[F_{S}]$. We can write the optimization problem for calculating such a budget as
@@
$C_{f}\geq 0$ $\forall f\in[F_{S}].$
In the above setting, we only obtain one single meta-solution (budget capacities for each category) to cover all scenarios. Note that the specific realization of the solution still has some leeway, as the specific selection of items is not prescribed by the meta-solution: it only tells us to use a budget of $C_{f}$ for items of category $f$ but does not prescribe which exact items to pick. An even more flexible alternative would be to have multiple possible meta-solutions (though fewer than the number of scenarios) and choose one among them for any given scenario. We index this set of meta-solutions by $k\in[K]$. We can then write the problem as
@@
s.t. $\sum_{f\in[F_{S}]}C_{f}^{k}\leq C$ $\forall k\in[K]$
$\sum_{i\in I_{f}}w_{i}x_{ji}\leq C_{f}^{k}+M(1-\ell^{k}_{j})$ $\forall f\in[F_{S}],\ j\in[N],\ \forall k\in[K]$
@@
$C^{k}_{f}\geq 0$ $\forall f\in[F_{S}],\ k\in[K]\,$,
where the goal is to find a function $f(\bm{c}_{j},\bm{\theta})$ that always takes the form of a decision tree. The splits $\bm{\theta}$ must be optimized so that the function maps instance features (the profit vector) to a meta-solution through this decision tree while maximizing the average profit. Note that this formulation is based on the one presented in (Bertsimas and Dunn, 2017), but other formulations exist as well (Aghaei et al., 2024).
We can understand the final knapsack problem in terms of the elements introduced in Section 2.1. Here, the set $\mathcal{X}$ corresponds to the budget constraint of the knapsack problem, which in this example is not affected by the instance features. However, the constraints forming the set are modified to model the meta-solutions. Each instance feature $\bm{\phi}_{I}^{j}$ reflects a different cost vector for the objective. The knapsack problem, as modeled, is equivalent to assigning equal probabilities to all scenarios. The solution features $\bm{\phi}_{S}$, thus correspond to a meta-solution, which is equivalent to a capacity vector $(C_{1}^{k},\dots,C_{F_{S}}^{k})$ with $\Phi_{S}\subseteq\mathbb{R}^{F_{S}}$. Given a meta-solution and instance features, the function $C(\bm{\phi}_{I},\bm{\phi}_{S})$ from our framework calculates the optimal objective value obtained from selecting the best singular solution obeying the meta-solution on that instance. The set $A$ is the space of all decision trees that map instance features to a meta-solution. We parameterize this set by $\bm{\theta}$. Finally, $\mu_{\bm{p}}$ represents how the objective values of different instances are aggregated, which is by summing them together, which is equivalent to the Laplace criteria.
@@
Assume we are given a graph $G=(V,E)$ representing a road network and nodes $s,t\in V$ representing a starting point and a destination. Furthermore, let us assume that traffic follows a pattern as shown in Figure 2. We assume that there are four main traffic scenarios that can occur at different times of the day throughout the week. We want to provide a comprehensible optimization rule for selecting a route, depending on the time of day and the day of the week.
Figure 2: Traffic scenarios assumed in the shortest path example. The day of the week is depicted on the horizontal and the time of day on the vertical axis.
To make our example more specific, we are using the graph from Goerigk et al. (2025), representing Chicago’s street network, which we have divided into eleven districts (see gray boxes in Figure 3).
@@
The edges are colored if the travel speed on them is slower than usual. The colors range from yellow (slightly reduced) to dark red (significantly slower) and provide information about the intensity of the speed reduction.
For the task of traveling from the northwest to the southeast of the city, let us first consider solutions given by the solution-based interpretable optimization rules.
@@
Figure 4: Optimization rule obtained from using the solution-based framework for interpretable surrogates.
Figure 4 illustrates them through a binary decision tree with a maximum depth of two. In each of the four scenarios identified by the decision tree’s splits, a specific $s$-$t$ path must be determined.
However, this approach may be too restrictive and diverges from how directions are typically explained in real life, where one would emphasize prominent landmarks, districts, streets, or intersections along the route, rather than detailing every road segment to be traveled. Therefore, we want to provide an optimization rule that uses solution features. In Figure 5, another binary decision tree is provided which functions as an optimization rule. However, instead of providing $s$-$t$ paths in the leaves, we now provide meta-solutions (shown in Figure 6), i.e., we provide features that any solution should possess in the given scenario. In this specific case, we select as solution features the sequence of city districts that one should traverse to reach the destination.
@@
(b) Meta-solution 2 and 4.
(c) Meta-solution 3.
Figure 6: Meta-solutions for the new framework. Districts shaded in gray have to be visited as per the order indicated by the bold gray arrows. Unshaded districts may not be used.
This way, we obtain a more comprehensible specification (specify a few features, rather than a lot of edges) and give the user more leeway in the implementation. In particular, any solution that follows the specified path through the districts is suitable.
This example also highlights the importance of adequately choosing solution features: E.g., when arriving as a tourist at O’Hare International Airport and searching for the most convenient way to your hotel, a specific route would probably be more helpful than providing district names the tourist is unfamiliar with. In contrast, a courier likely has this kind of knowledge and providing landmarks would suffice in order to guide the route. Hence, it is evident that the solution features have to be carefully matched to the capabilities and expert knowledge of the user.